i need a solution for these provlems. plz~~
3) The vibrational energy of a carbon monoxide, co
Answer #1

3) Ans b

Vibrational energy(E) = potential energy = 1/2kx^{2}

E=1/2kx^{2} where k=force constant, x = classical amplitude

  x = \sqrt{2E/k} = \sqrt{2*3.42*10^{-20}/2170} = 5.61 * 10^{-12}m = 5.61 pm

4) Ans a

IE = RZ^{2}[1/n_{1}^{2}-1/n_{2}^{2}]

where IE = ionization energy, R: rydberg constant, n1= 1 and n2= infinity, Z = atomic number Be3+ = 4

IE = 109731*4*4 [1-0]= 1755696 cm-1/8065 = 218 ev

5) Ans d

selection rule for orbital angular momentum

  \Delta l =+/- 1

for 5d , l = 2

possible transitions to l = 1 i.e p and l = 3 i.e f

therefor np and nf